# PerLA  [Uncaptioned image] : Perceptive 3D language assistant

### v1

@@ 5 Conclusions
Limitations & Future Work PerLA primarily focuses on enhancing performance using point cloud input, but integrating it with optimized modules, such as token merging, presents a promising direction to extend its capabilities. Although PerLA shows strong performance on standard benchmarks, future work could explore its robustness and generalizability in outdoor environments and more complex 3D scenarios, broadening its applicability in diverse real-world settings.
@@ B.2 Multimodal prompts
Scene-aware visual prompt encoder.
The visual prompt encoder aims to process visual prompts into representations that are then easier to be processed by the LLM. In addition to prior work which applies positional encoding followed by an MLP layer to process the visual prompts (as described in Eq. 3 of [9]), we further enhance this approach by augmenting the positional encoding with the detail-enriched global scene representations ^$\hat{\mathcal{F}}^{g}$, obtained by our perceptive scene encoder. Therefore, the visual prompt encoder is more aligned to the scene representation, helping to improve the model performance as empirically proved in Tab. I.
@@
Textual prompt encoder.
Textual prompts provide task-specific instructions to 3DLAs. For 3D dense captioning, we instruct the model to perform one of two tasks: “describe" or “describe and localize" the object, while for 3D question answering, we use textual instructions that ask the model to either “answer" or “answer and localize the related objects." Specifically, we encode the input text prompt $\mathcal{I}^{t}$ using a transformer architecture inspired by BLIP-2 [28, 9]. This transformer is initialized with a pre-trained BERT model to handle word and positional embeddings, producing text representations $\mathcal{F}^{e}\in\mathbb{R}^{T\times d_{e}}$.
@@ Appendix D Dataset Details
Nr3D dataset [1] is a benchmark for 3D object localization tasks in natural language, built on the ScanNet [14] dataset. It contains 41,503 unique natural language descriptions referring to 5,578 objects across 707 ScanNet scenes. Each description is designed to unambiguously identify a target object in the context of its surrounding scene, incorporating spatial relationships and object attributes such as color, shape, and size. On average, each object is associated with 7.4 descriptions, providing comprehensive linguistic diversity. The dataset focuses on common indoor objects, making it suitable for evaluating fine-grained understanding of object attributes and spatial reasoning in 3D scenes.
@@ E.1 Details with increasing tokens (LL3DA)
In the main paper (Tab. 3), we introduce a variant of LL3DA, which combines both global and local information by leveraging query tokens generated from local and global regions. LL3DA serves as a baseline of enriching global context with local details by extending the number of query tokens. LL3DA first extracts 3D representations from different partitions of the scene. It then applies Farthest Point Sampling (FPS) to select 1,024 points, along with their corresponding point-level representations, from the union of these partitions. These sampled representations are processed through the multimodal adapter to produce 32 local tokens. Then, we can generate an additional 32 tokens from the point cloud of the entire scene. We finally obtain a total of 64 tokens by concatenating such local and global tokens and processed through a self-attention layer. The self-attention mechanism enables interaction and information exchange between the 64 tokens, enhancing the representation of both global and local representations. We evaluate the performance of LL3DA on the ScanQA validation dataset, in comparison to original LL3DA and our PerLA. As demonstrated in Tab. H, increasing the number of tokens allows LL3DA to achieve a modest performance improvement, as expected. However, its effectiveness remains limited compared to PerLA. This is because PerLA is specifically designed to capture both global context and local details during the scene encoding phase. Furthermore, it is trained with carefully crafted loss functions that ensure stable and efficient learning.
@@ E.2 Scene-aware visual prompt encoder
E.3 Performance for two stage training
Table J: Performance for Two-Stage Training. The first three rows report the performance of PerLA trained from scratch as task-specific experts on their respective training datasets. The subsequent three rows present the results of models fine-tuned on each dataset using weights initialized from the generalist model trained across all task-specific datasets. The final row evaluates the generalist model’s performance without fine-tuning. ScanRefer [6] and Nr3D [1] are evaluated for dense captioning, while ScanQA [2] is evaluated for question answering. The results demonstrate the effectiveness of our generalist model in multi-task scenarios and its strong performance after fine-tuning.
@@ E.3 Performance for two stage training
Table K: Quantitative Comparisons on Scene Description, Embodied Dialogue, and Embodied Planning on the ScanNet part of 3D-LLM [21] with a beam size = 4 for beam search.
@@
We evaluate the performance of PerLA under this two-stage training paradigm, as summarized in Tab. J, together with several baseline methods. The first three rows in Tab. J display the performance of models trained from scratch as task-specific experts. The next three rows show the results of models fine-tuned on individual tasks, initialized from the generalist model’s weights. The final row reports the performance of the generalist model without fine-tuning, where a single set of weights is used to handle all tasks.
@@
E.4 More results on Scene Description, Embodied Dialogue and Embodied Planning
@@
As shown in Tab. K, for scene description, PerLA outperforms all baselines, including LL3DA, across all metrics. Notably, it achieves substantial improvements in CiDER (+3.07) and METEOR (+1.54) compared to LL3DA.

### v2

@@ 5 Conclusions
Limitations & Future Work. PerLA primarily focuses on enhancing performance using point cloud input, but integrating it with optimized modules, such as token merging, presents a promising direction to extend its capabilities. Although PerLA shows strong performance on standard benchmarks, future work could explore its robustness and generalizability in more complex 3D scenarios, broadening its applicability in diverse real-world settings.
Acknowledgement. This work was supported by PNRR FAIR - Future AI Research (PE00000013) and ICSC National Research Centre for HPC, Big Data and Quantum Computing (CN00000013), funded by NextGeneration EU.
@@ B.2 Multimodal prompts
Scene-aware visual prompt encoder. The visual prompt encoder aims to process visual prompts into representations that are then easier to be processed by the LLM. In addition to prior work which applies positional encoding followed by an MLP layer to process the visual prompts (as described in Eq. 3 of [9]), we further enhance this approach by augmenting the positional encoding with the detail-enriched global scene representations ^$\hat{\mathcal{F}}^{g}$, obtained by our perceptive scene encoder. Therefore, the visual prompt encoder is more aligned to the scene representation, helping to improve the model performance as empirically proved in Tab. I.
@@
Textual prompt encoder. Textual prompts provide task-specific instructions to 3DLAs. For 3D dense captioning, we instruct the model to perform one of two tasks: “describe" or “describe and localize" the object, while for 3D question answering, we use textual instructions that ask the model to either “answer" or “answer and localize the related objects." Specifically, we encode the input text prompt $\mathcal{I}^{t}$ using a transformer architecture inspired by BLIP-2 [31, 9]. This transformer is initialized with a pre-trained BERT model to handle word and positional embeddings, producing text representations $\mathcal{F}^{e}\in\mathbb{R}^{T\times d_{e}}$.
@@ C.2 Computation efficiency
Hilbert-based nearest-neighbor search. Our $k$-NN search search is constrained by geometric segments. We first serialize the union of the global and local point clouds $\mathcal{P}^{g}\cup\mathcal{P}^{l}$, using a Hilbert curve ordering. This ordering maps each point’s multi-dimensional coordinates to a one-dimensional Hilbert bit $d$, thereby preserving spatial locality. Next, we apply geometric partitioning [27] on the original point clouds to generate geometric labels $\mathcal{Y}^{g}$ and $\mathcal{Y}^{l}$. For each label, we compute its center (i.e., the mean coordinate of all points in that label) and incorporate this center information into the serialized index as high-order bits via bit shifting. In practice, this is achieved by computing a combined metric for each point:
[…]
where label_offset is chosen to be larger than the maximum possible Hilbert distance. Once the union is sorted per batch using this combined index, we perform an approximate $k$-NN search.
@@ Appendix D Dataset Details
Nr3D dataset [1]
is a benchmark for 3D object localization tasks in natural language, built on the ScanNet [14] dataset. It contains 41,503 unique natural language descriptions referring to 5,578 objects across 707 ScanNet scenes. Each description is designed to unambiguously identify a target object in the context of its surrounding scene, incorporating spatial relationships and object attributes such as color, shape, and size. On average, each object is associated with 7.4 descriptions, providing comprehensive linguistic diversity. The dataset focuses on common indoor objects, making it suitable for evaluating fine-grained understanding of object attributes and spatial reasoning in 3D scenes.
@@ E.1 Details with increasing tokens (LL3DA)
In the main paper (Tab. 3), we introduce a variant of LL3DA, which combines both global and local information by leveraging query tokens generated from local and global regions. LL3DA serves as a baseline of enriching global context with local details by extending the number of query tokens. LL3DA first extracts 3D representations from different partitions of the scene. It then applies Farthest Point Sampling (FPS) to select 1,024 points, along with their corresponding point-level representations, from the union of these partitions. These sampled representations are processed through the multimodal adapter to produce 32 local tokens. Then, we can generate an additional 32 tokens (global tokens) from the point cloud of the entire scene. We finally obtain a total of 64 tokens by concatenating such local and global tokens and processed through a self-attention layer, i.e., $\mathcal{Q}^{64\times 768}=f_{att}\left(\mbox{cat}[\mbox{MMA}(\mathcal{F}^{g},\mathcal{I}^{t}),\mbox{MMA}(\mathcal{F}^{l},\mathcal{I}^{t})]\right)$ (where the number of tokens increases from 32 to 64). The self-attention mechanism enables interaction and information exchange between the 64 tokens, enhancing the representation of both global and local representations. We evaluate the performance of LL3DA on the ScanQA validation dataset, comparing it with the original LL3DA and our PerLA. As shown in Tab. H, increasing the number of tokens enables LL3DA to achieve a modest performance gain over the reproduced LL3DA (repr.), as expected. However, its effectiveness remains limited compared to PerLA. This is because PerLA is specifically designed to capture both global context and local details during the scene encoding phase. Furthermore, it is trained with carefully crafted loss functions that ensure stable and efficient learning. Our findings align with Idefics2 [28]. Idefics2 shows that reducing the number of visual tokens through attention-based pooling significantly enhances computational efficiency during training and inference while improving performance on downstream tasks.
@@ E.2 Scene-aware visual prompt encoder
Table J: Ablation study on the impact of the numbers of learnable query tokens on the ScanQA [2] validation dataset.
Method 32 64 96 128
[…]
LL3DA 73.2 12.8 14.9 36.0 73.7 13.6 15.2 35.5 74.2 13.8 15.1 36.1 74.6 13.5 15.2 36.0
PerLA 74.4 13.7 15.8 36.5 75.1 14.2 15.9 36.6 76.4 14.3 16.3 40.0 77.0 14.3 16.4 38.2
E.3 Increase number of tokens
To evaluate the impact of the number of learnable query tokens, we experimented with various numbers of learnable query tokens, from 32 to 128, in the Q-Former (different from LL3DA). We trained each model from scratch on the ScanQA training set and evaluated its performance on the validation set for both LL3DA and PerLA. We observed that increasing the number of tokens to 96 and 128 led to loss divergence (to infinity) in both LL3DA and PerLA. To address this instability, we incorporated mirror descent-based regularization [58]. Tab. J presents the results. Both LL3DA and PerLA achieve better performance with increasing query tokens, where the performance gain tends to saturate from 96 to 128 tokens. Moreover, our proposed method PerLA consistently outperforms LL3DA across all token configurations and evaluation metrics.
E.4 Performance for two stage training
Table K: Performance for Two-Stage Training. The first three rows report the performance of PerLA trained from scratch as task-specific experts on their respective training datasets. The subsequent three rows present the results of models fine-tuned on each dataset using weights initialized from the generalist model trained across all task-specific datasets. The final row evaluates the generalist model’s performance without fine-tuning. ScanRefer [6] and Nr3D [1] are evaluated for dense captioning, while ScanQA [2] is evaluated for question answering. The results demonstrate the effectiveness of our generalist model in multi-task scenarios and its strong performance after fine-tuning.
@@ E.4 Performance for two stage training
Table L: Quantitative Comparisons on Scene Description, Embodied Dialogue, and Embodied Planning on the ScanNet part of 3D-LLM [21] with a beam size = 4 for beam search.
@@
We evaluate the performance of PerLA under this two-stage training paradigm, as summarized in Tab. K, together with several baseline methods. The first three rows in Tab. K display the performance of models trained from scratch as task-specific experts. The next three rows show the results of models fine-tuned on individual tasks, initialized from the generalist model’s weights. The final row reports the performance of the generalist model without fine-tuning, where a single set of weights is used to handle all tasks.
@@
E.5 More results on Scene Description, Embodied Dialogue and Embodied Planning
@@
As shown in Tab. L, for scene description, PerLA outperforms all baselines, including LL3DA, across all metrics. Notably, it achieves substantial improvements in CiDER (+3.07) and METEOR (+1.54) compared to LL3DA.